The admissions director of an engineering college has $400,000 in scholarships each year from an endowment to offer to high-aPlease use excel to answer the question.
Answer #1

A B B С 1 2 3 Number of Offers Acceptance Percentage Scholarship Amount Accepted Scholarship Available Scholarship Extra FundWe have Number of offers = 50

Accepted scholarships = 60% of 50 = 30

Total funds needed= 30* 25,000 = 750,000

Funds available = 400,000

Extra funds needed = 750,000 - 400,000 = 350,000